OpenStudy (ilovelife):

HELPPPPP ! What is the solution set of |2x – 3| – 8 = –1?

OpenStudy (anonymous):

I think it is 5 or -5

OpenStudy (anonymous):

kinda rusty with these though

OpenStudy (barrycarter):

Hint: add 8 to both sides

OpenStudy (ilovelife):

the options are ... 7 or –7 5 or –2 5 or 2 No solution

OpenStudy (barrycarter):

Hint: if |x| = something, then x is either that something or negative that something

OpenStudy (campbell_st):

add 8 to both sides |2x - 3| = 7 then you have 2 equations to solve 2x - 3 = 7 and 2x - 3 = -7

OpenStudy (ilovelife):

i got -5 and 2 but those aren't in the choices.

jimthompson5910 (jim_thompson5910):

I'll solve the first equation and (hopefully) you can use this to solve the second 2x - 3 = 7 2x - 3+3 = 7+3 2x = 10 2x/2 = 10/2 x = 5

jimthompson5910 (jim_thompson5910):

So I'm not sure how you got -5 instead of 5

OpenStudy (ilovelife):

i did +3 by accident -___- lol. THANKS GUYS !

OpenStudy (ilovelife):

big help(: